# Asymmetrical temporal dynamics of topological edge modes in Su-Schrieffer-Heeger lattice with Kerr nonlinearity

### v1

@@ VI Conclusions
In summary, we have numerically demonstrated the optical bistability and various types of oscillations in a 1D SSH model composed of ring resonators with Kerr nonlinearity. When the nonlinear terms are introduced in the Lugiato-Lefever equation, the first ring’s CW and CCW mode intensities are symmetric until the pump intensity reaches a bifurcation point. Above the bifurcation point, the symmetry is spontaneously broken due to the splitting of the resonance frequencies of the two CW and CCW modes in the first ring resonator. For the high input intensity regime, we have observed oscillating phases including periodic and chaotic oscillations. The periodic oscillation phases can also be classified into two different phases where the trajectories are separate or identical in the phase space of the mode intensities. This emergence of the oscillating phases can be attributed to the coupling between the edge mode and bulk mode due to the large shift of resonance peaks of the edge mode.
We believe that our theoretical models and numerical results will provide valueable insight in understanding the complex dynamics in coupled nonlinear resonator systems with two chiral modes. Additionally, the various spatio-temporal dynamics could be applied to optical switching devices as well as the stability analysis of coupled lasers.
@@ Appendix A Derivation of the Lugiato-Lefever equation using the coupled mode theory
$\frac{d\tilde{a}}{dt}=[-\gamma-i\tilde{\Delta}]\tilde{a}+\gamma_{c}\tilde{s}.$ (15)
@@
Appendix B Site-dependence of Poincaré sections for nonlinear SSH model
@@
Figure A1: Poincaré sections of the maxima of oscillating coupled intensity as a function of detuning for a 1D SSH lattice ($M=7$) with the same parameters in Fig. 4. (a) For odd sites in the main text ($I_{s}=20,v=3,w=7,A=1,B=4$) with the cyan, green, blue, and red colors corresponding to the 1st, 3rd, 5th, and 7th ring resonators respectively. (b) For even sites with the red, green and blue colors corresponding to the 2nd, 4th, and 6th ring resonators, respectively.
Appendix C Poincaré sections of the maxima of oscillating coupled intensity as a function of B
Figure A2: Poincaré section of the maxima of oscillating coupled intensity as a function of $B$ for input intensity $I_{s}=20$, for the first ring resonator from the 1D SSH array. The Poincaré sections for coupling coefficients ($v=3$, $w=7$) and $\Delta=5.5$. Cyan shading indicates oscillations and yellow refers to optical bistability.
In Fig. A2 we scan $B$ from $1$ to $7$ to observe oscillations in coupled intensity for $A=1$, $I_{s}=20$, $v=3$ and $w=7$. Here, the Poincaré section can be divided into three regions: symmetric stable, asymmetric stable and asymmetric unstable states. We observe a series of bifurcations and oscillation windows for $B\geq 1.9$.

### v2

@@ VI Conclusions
In summary, we have numerically demonstrated the optical bistability and various types of oscillations in a 1D SSH model composed of ring resonators with Kerr nonlinearity. When the nonlinear terms are introduced in the Lugiato-Lefever equation, the first ring’s CW and CCW mode intensities are symmetric until the pump intensity reaches a bifurcation point. Above the bifurcation point, the symmetry is spontaneously broken due to the splitting of the resonance frequencies of the two CW and CCW modes in the first ring resonator. For the high intensity regime, we have observed oscillating phases including periodic and chaotic oscillations. We have classified the periodic oscillation phases into two different phases where the trajectories are separate or identical in the phase space of the mode intensities. This emergence of the oscillating phases can be attributed to the coupling between the edge mode and bulk mode due to the large shift of resonance peaks of the edge mode.
We believe that our theoretical model and numerical results will provide valuable insight in understanding the complex dynamics in coupled nonlinear resonator systems with two chiral modes. In practice, the asymmetrical dynamic modes can be demonstrated experimentally using an optical ring resonator array, such as, a silicon microring resonator array [42]. In particular, the bifurcation states can be a building block for optical memory and switching devices allowing for storing and maintaining information by controlling the pump intensities depending on the direction of light circulation [43, 44]. Additionally, the various spatio-temporal dynamics could be applied to the stability analysis of coupled lasers.
@@ Appendix A Derivation of the Lugiato-Lefever equation using the coupled mode theory
$\frac{d\tilde{a}}{dt}=(-\gamma-i\tilde{\Delta})\tilde{a}+\gamma_{c}\tilde{s}.$ (20)
@@
Appendix B Sublattice symmetry and CW-CCW symmetry
In the linear regime, i.e., when $I_{s}\simeq 0$, our SSH model (Fig. 1(b)) has two distinctive symmetries called sublattice symmetry and CW-CCW symmetry. While both of them are called chiral symmetry in the literature, we use the terms sublattice symmetry and CW-CCW symmetry to avoid confusion. The Hamiltonian of the SSH model can be written in the following block matrix form:
$\mathbf{H}=\begin{pmatrix}\mathbf{H}_{cw}&\mbox{\Large 0}\\ \mbox{\Large 0}&\mathbf{H}_{ccw}\par\end{pmatrix},$ (21)
where $\mathbf{H}_{cw}$ ($\mathbf{H}_{ccw}$) is the SSH Hamiltonian for CW (CCW) modes with the size of $M\times M$. They can be written as:
$\mathbf{H}_{cw}=\begin{pmatrix}\mu_{1,a}^{cw}&v&0&0&\ldots\\ v&\mu_{1,b}^{cw}&w&0&\ldots\\ 0&w&\ddots&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots\end{pmatrix},$ (22)
and
$\mathbf{H}_{ccw}=\begin{pmatrix}\mu_{1,a}^{ccw}&v&0&0&\ldots\\ v&\mu_{1,b}^{ccw}&w&0&\ldots\\ 0&w&\ddots&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots\end{pmatrix},$ (23)
respectively. Here, the diagonal terms, corresponding to the onsite energy in a tight-binding description, represent the shift of resonance frequencies:
$\mu_{n,a}^{cw}=-\eta\Delta+A|\tilde{a}_{n,+}|^{2}+B|\tilde{a}_{n,-}|^{2},$ (24)
and
$\mu_{n,a}^{ccw}=-\eta\Delta+A|\tilde{a}_{n,-}|^{2}+B|\tilde{a}_{n,+}|^{2}.$ (25)
$\mu_{n,b}^{cw}$ and $\mu_{n,b}^{ccw}$ are defined similary.
First, let us check the sublattice symmetry of the nonlinear SSH Hamiltonian. The sublattice symmetry operator $\mathbf{\Gamma}$ for the subspace with the size of $M\times M$ is given by :
$\mathbf{\Gamma}=\begin{pmatrix}-1&0&0&\ldots\\ 0&1&0&\ldots\\ 0&0&-1&\ddots\\ \vdots&\vdots&\ddots&\ddots\end{pmatrix}.$ (26)
If we apply $\mathbf{\Gamma}$ to $\mathbf{H}_{cw}$ ($\mathbf{H}_{ccw}$), we obtain
$\mathbf{\Gamma}\mathbf{H}_{cw}\mathbf{\Gamma}^{\dagger}=\begin{pmatrix}\mu_{1,a}^{cw}&-v&0&0&\ldots\\ -v&\mu_{1,b}^{cw}&-w&0&\ldots\\ 0&-w&\ddots&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots\end{pmatrix},$ (27)
and
$\mathbf{\Gamma}\mathbf{H}_{ccw}\mathbf{\Gamma}^{\dagger}=\begin{pmatrix}\mu_{1,a}^{ccw}&-v&0&0&\ldots\\ -v&\mu_{1,b}^{ccw}&-w&0&\ldots\\ 0&-w&\ddots&\ddots&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots\end{pmatrix}.$ (28)
This means that the sublattice symmetry is broken in the strong nonlinear regime because the diagonal terms of $\mathbf{H}_{cw}$ and $\mathbf{H}_{ccw}$ do not vanish making $\mathbf{\Gamma}\mathbf{H}\mathbf{\Gamma}^{\dagger}\neq-\mathbf{H}$. In the weak nonlinearity regime, however, the diagonal terms approximately equal zero, satisfying $\mathbf{\Gamma}\mathbf{H}\mathbf{\Gamma}^{\dagger}=-\mathbf{H}$, the sublattice symmetry is recovered. Here, we assume the detuning is zero.
Next, we check the CW-CCW symmetry. Let us define a block mirror symmetry operator matrix $\mathbf{P}$ as:
$\mathbf{P}=\begin{pmatrix}\mbox{\Large 0}&\mathbb{I}_{M}\\ \mathbb{I}_{M}&\mbox{\Large 0}\end{pmatrix},$ (29)
where $\mathbb{I}_{M}$ is an identity matrix with size of $M\times M$. Then, we apply $\mathbf{P}$ to the full Hamiltonian as below.
$\mathbf{P}\mathbf{H}\mathbf{P}^{-1}=\begin{pmatrix}\mathbf{H}_{ccw}&\mbox{\Large 0}\\ \mbox{\Large 0}&\mathbf{H}_{cw}\end{pmatrix}.$ (30)
From 22, 23, one can notice that $\mathbf{P}\mathbf{H}\mathbf{P}^{-1}=\mathbf{H}$ is satisfied only when $\tilde{a}_{n,-}=\tilde{a}_{n,+}$ which corresponds to the symmetrical dynamic modes in the weak nonlinear regime. For strong nonlinear regime, we obtain $\mathbf{P}\mathbf{H}\mathbf{P}^{-1}\neq\mathbf{H}$ giving rise to asymmetrical dynamics.
Appendix C Site-dependence of Poincaré sections for nonlinear SSH model
@@
Figure A1: Poincaré sections of the maxima of oscillating coupled intensity as a function of detuning for a 1D SSH lattice ($M=7$) with the same parameters in Fig. 4(a). (a) For odd sites in the main text ($I_{s}=40,v=3,w=7,A=1,B=4,\gamma_{c}=1$) with the cyan, green, blue, and red colors corresponding to the 1st, 3rd, 5th, and 7th ring resonators respectively. (b) For even sites with the red, green and blue colors corresponding to the 2nd, 4th, and 6th ring resonators, respectively.